OpenStudy (anonymous):

Write the expression in radical form: (x^3y^3)^1/6

OpenStudy (anonymous):

\[\left( x^3y^3 \right)^{1/6}\]

OpenStudy (anonymous):

Example. \[(x^5y^5)^\frac{ 1 }{ 4 } = \sqrt[4]{x^5y^5}\] Hope that helps

OpenStudy (anonymous):

Our teacher simplifies first and gets \[\sqrt{xy}\] I just don't know how he got that answer?

OpenStudy (nikato):